# How Temporal Unrolling Supports Neural Physics Simulators

### v1

@@ 3.2 Unrolling dynamical systems at training time
Unrolling describes the process of auto-regressively evolving the learned system in a training iteration. It can be formalized as
$\mathbf{u}^{t+s}=g_{\theta}^{s}(\mathbf{u}^{t}),$ (2)
where $g^{s}$ represents the recurrent application of $s$ steps of either correction, for which $g_{\theta}(\mathbf{u})=f_{\theta}(\mathcal{S}(\mathbf{u}))$, or prediction, where $g_{\theta}(\mathbf{u})=f_{\theta}(\mathbf{u})$. Similarly to the target dynamics, the learned system also has an attractor $A_{g_{\theta}}$ defining its solution space. When unrolling a learning trajectory, multiple frames $[\mathbf{u}^{t+1},\mathbf{u}^{t+2},\dots,\mathbf{u}^{t+m}$] are generated, where $m$ denotes the number of unrolled time steps. A simple one-step training is obtained by setting $m=1$. The objective of the neural network is to fit the PDE dynamics in equation 1. The network models are trained to represent the underlying PDE by fitting a dataset of ground truth trajectories $\tilde{U}=[\mathbf{\tilde{u}}^{0}\dots\mathbf{\tilde{u}}^{N}]$. This dataset contains samples from the target attractor $A_{\mathcal{F}}$, which represents a subset of the phase space of $\mathbf{\tilde{u}}$ restricted to physical solutions. Nevertheless, due to imperfections in the learned solution, the learned dynamics (2) will differ from the ground truth at training and inference time. Thus, the learned attractor $A_{g_{\theta}}\neq A_{\mathcal{F}}$.
Crucially, training without unrolling does not guarantee that the states $\mathbf{u}$ observed during training sufficiently represent $A_{g_{\theta}}$. In other words,
@@
$\forall\mathbf{\tilde{u}}^{i}\in A_{\mathcal{F}},A_{g_{\theta}}:\lim_{N\rightarrow\infty}\{g_{\theta}(\mathbf{\tilde{u}}^{0}),g_{\theta}(\mathbf{\tilde{u}}^{1}),...,g_{\theta}(\mathbf{\tilde{u}}^{N})\}\neq A_{g_{\theta}}.$ (3)
@@
This means that we are not guaranteed to explore the attractor of the learned system $A_{g_{\theta}}$ when only observing states based on one discrete evolution of $g_{\theta}$, regardless of the dataset size $N$. A common data-augmentation approach is to perturb the ground truth states in $\tilde{U}$ with noise, in the hope that the perturbed states fit $A_{g_{\theta}}$ more closely [51]. While this approach was reported to be successful, it still does not provide a guarantee that the observed states are physical under $g_{\theta}$.
@@
Let us now suppose we unroll $m$ steps with $g_{\theta}^{s}(\mathbf{\tilde{u}})$. Based on the definition of the attractor $A_{g_{\theta}}$ we can state that
$\lim_{m\rightarrow\infty}\{g_{\theta}^{0}(\mathbf{\tilde{u}}^{i}),g_{\theta}^{1}(\mathbf{\tilde{u}}^{i}),...,g_{\theta}^{m}(\mathbf{\tilde{u}}^{i})\}=A_{g_{\theta}},$ (4)
@@
for any $\mathbf{\tilde{u}}^{i}$ in the vicinity of $A_{g_{\theta}}$. While the horizon $m$ is practically never chosen long enough to expose $A_{g_{\theta}}$ with a single initial condition, a sufficiently large dataset ensures that $A_{g_{\theta}}$ is approached from multiple angles. In contrast to training wihtout unrolling, enabling it thus exposes the inference attractor $A_{g_{\theta}}$ at training time for sufficiently large datasets and horizons $m$. Precisely the difference between the observed training set (sampled from $A_{\mathcal{F}}$) and the inference attractor $A_{g_{\theta}}$ is commonly known as data shift in machine learning. The dark red curve in Figure 1 shows how this data shift is thus reduced by choosing larger $m$.
@@ 3.3 Training neural networks to solve PDEs:
The general training procedure is given as
$\theta^{*}=\mathrm{arg}\min_{\theta}\bigg{[}\sum_{i=1}^{N-m}\sum_{s={1}}^{m}\mathcal{L}_{2}\big{(}\tilde{\mathbf{u}}^{i+s\tau},g_{\theta}^{s}(\mathbf{u}^{i})\big{)}\bigg{]},$ (5)
where $\tau$ accounts for the relative timestep between ground truth and predicted trajectories. For one learning iteration, a loss value is accumulated over an unrolled trajectory. We can write this total loss over the unrolled trajectory as
$\mathcal{L}=\sum_{i=1}^{s}\mathcal{L}_{2}(\tilde{\mathbf{u}}^{i+\tau s},g^{s}_{\theta}(\mathbf{u}^{i}))=\sum_{i=1}^{s}\mathcal{L}^{s},$ (6)
where $\mathcal{L}^{s}$ represents the loss evaluated after a step. The simplest realization of the training procedure (5) utilizes the forward process with $m=1$, i.e. no unrolling. We refer to this variant without unrolling as one-step (ONE) in the following. This approach is common practice in literature and the resulting training procedure is easily realized with pre-computed data. In contrast, the full backpropagation through an unrolled chain with $m>1$ requires a differentiable solver $\mathcal{S}$ for the correction setup. The differentiable setup can calculate the full optimization gradients by propagating gradients through a simulation step. The gradients are thus evaluated as
@@
We refer to this fully differentiable setup as with-gradient (WIG). The gradient flow through a simulator step $g$ is caluclated as $\frac{\partial g_{\theta}^{s+1}}{\partial g_{\theta}^{s}}=\frac{\partial f_{\theta}^{s}}{\partial\mathcal{S}^{s}}\frac{\partial\mathcal{S}^{s}}{\partial g_{\theta}^{s}}$ for correction setups and as $\frac{\partial g_{\theta}^{s+1}}{\partial g_{\theta}^{s}}=\frac{\partial f_{\theta}^{s}}{\partial g_{\theta}^{s}}$ for prediction. To test the effect of differentiability, we introduce a second strategy for propagating the gradients $\frac{\partial\mathcal{L}^{s}}{\partial\theta}$. If no differentiable solver is available, optimization gradients can only propagate to the network application, not through the solver. The gradients are thus evaluated as
@@
This setup is referred to as no-gradient (NOG). Here, $\frac{\partial g_{\theta}^{s+1}}{\partial g_{\theta}^{s}}:=0$ holds for correction and prediction setups. Thus, a simulator step in the NOG setup is not differentiable and yields no gradients. Most existing code bases in engineering and science are not fully differentiable. Consequentially, this NOG setup is particularly relevant, as it could be implemented using existing traditional numerical solvers.
@@
Figure 2 contains a visualization including the backpropagation flow through an unrolled chain. The gradient calculations for our three setups are further denoted in table 1. Further detail of the gradient calculations in correction and prediction setups is found in A.
@@ 3.4 Gradient behavior in unrolled systems
Integrating a property of chaotic dynamics derived by [42] we observe the Jacobians $\mathbf{J_{s}}=\frac{\partial g_{\theta}^{s}}{\partial g_{\theta}^{s-1}}$ have eigenvalues larger than 1 in the geometric mean. Thus,
$\bigg{|}\bigg{|}\prod_{b=1}^{s-B}\frac{\partial g_{\theta}^{s-b}}{\partial g_{\theta}^{s-b-1}}\bigg{|}\bigg{|}>1$ (10)
holds in this case. This means that the gradient inaccuracy increases with the unrolling length. We can also observe that the NOG gradient inaccuracy grows with $\propto m^{2}$, while the NOG gradients only linearly depend on $m$. As a consequence, the gradients computed in the NOG system diverge from the true (WIG) gradients for increasing $m$, as shown with the blue line in Figure 1. The gradient approximations used in NOG thus do not accurately match the loss used in training. This hinders the network optimization.
Let us finally consider the full gradients of the WIG setup. We can use Theorem 2 from [42], which states that
@@
for almost all points on $A_{g_{\theta}}$. The gradients of the WIG setup explode exponentially for long horizons $m$, but less quickly than those of NOG, as indicated by the orange line below the blue one in Figure 1. As a direct consequence, the training of the WIG setup becomes unstable for chaotic systems when $m$ grows to infinity.
We can summarize the above findings as follows. Unrolling training trajectories for data-driven learning of chaotic systems reduces the data-shift, as the observed training samples converge to the learned attractor. At the same time, long unrollings lead to unfavorable gradients for both NOG and WIG setups. This means that intermediate ranges of unrolling horizons can exist, where the benefits of reducing the data-shift outweigh instabilities in the gradient.
As part of our numerical experiments, we expose these unrolling horizons and quantify the expected benefits of training with NOG and WIG modalities.
Figure 2: Left: Illustration of unrolling for a horizon $m=2$ for correction and prediction chains, including gradient flow in the backward pass for non-differentiable (blue, NOG) and differentiable (orange, WIG) setups, note how the gradients do not flow through a time step for NOG; Right: Contributions of networks and numerical solvers in prediction and correction paradigms
@@ 4 Physical Systems and Architectures
Four physical systems were used for our learning tests. All systems are parameterized to exhibit varying behavior, and each test set contains unseen values inside and outside the range of the training data set.

### v2

@@ 2.3 Unrolling dynamical systems at training time
In practice, Neural Simulators rarely satisfy Equation (3) exactly and rather form an approximation of the ground truth. Over long auto-regressive rollouts at inference time these small-scale errors accumulate and lead to a substantial shift in the solution, which can cause instabilities and ultimately makes the solution diverge. The data shift between snapshots observed at training time and those observed in inference is commonly identified as the cause of these instabilities [72]. When simply training on one-step trajectories as in Equation (3), only data sampled from the ground truth attractor $A_{\mathcal{F}}$ is observed, as only data from the ground truth dataset $\tilde{\mathbf{u}}\in\tilde{U}$ is supplied. The same model might then used autoregressively at inference time with $\mathbf{u}^{i+1}=g_{\theta}(\mathbf{u}^{i})$. An inference trajectory is then constructed as $[\mathbf{u}^{i+1},g_{\theta}(\mathbf{u}^{i+1}),\dots]$ Crucially, we supply the model with self-generated data $\mathbf{u}_{i}$ for the first time when computing this inference trajectory. Based on the fact that learned simulators never perfectly reproduce the ground truth, this can lead to error accumulation and unwanted behavior in the trajectories. This means, that the attractor of the learned dynamics $A_{g_{\theta}}$ includes additional states in such cases. These states are exposed by the auto-regressive inference rollout which feeds unseen and fundamentally different data to the network. Thus, the learned attractor $A_{g_{\theta}}\neq A_{\mathcal{F}}$.
@@
We study unrolling as a possible remedy for this problem. Unrolling describes the process of auto-regressively evolving the learned system in a training iteration. It can be formalized as
$\mathbf{u}^{t+s}=g_{\theta}^{s}(\tilde{\mathbf{u}}^{t}),$ (8)
where $g^{s}_{\theta}$ represents the recurrent application of $s$ steps of $g_{\theta}$. When unrolling at training time from a starting frame $\tilde{\mathbf{u}}^{t}$, multiple frames $[\mathbf{u}^{t+1},\mathbf{u}^{t+2},\dots,\mathbf{u}^{t+m}$] are generated within one training iteration, where $m$ denotes the unrolling horizon. A simple one-step training is obtained by setting $m=1$. However, one-step training without unrolling does not guarantee that the states $\mathbf{u}$ observed during training sufficiently represent $A_{g_{\theta}}$. In mathematical terms,
@@
$\forall\mathbf{\tilde{u}}^{t}\in A_{\mathcal{F}}\>\exists\>{g_{\theta}}:\lim_{N\rightarrow\infty}\{g_{\theta}(\mathbf{\tilde{u}}^{0}),g_{\theta}(\mathbf{\tilde{u}}^{1}),...,g_{\theta}(\mathbf{\tilde{u}}^{N})\}\neq A_{g_{\theta}}.$ (9)
@@
This means that we are not guaranteed to explore the attractor of the learned system $A_{g_{\theta}}$ when only those states are observed that are based on one discrete evolution of $g_{\theta}$, regardless of the dataset size $N$. Simple examples of this are networks sampled during training. A common data-augmentation approach is to perturb the ground truth states in $\tilde{U}$ with noise, in the hope that the perturbed states fit $A_{g_{\theta}}$ more closely [79]. While this approach was reported to be successful, it still does not guarantee that the observed states are from $A_{g_{\theta}}$, and thus, those augmented states are not necessarily closer to states observed during inference.
@@
Let us now suppose we unroll to the horizon $m$ with $g_{\theta}^{s}(\mathbf{\tilde{u}})$ and $s\in[1,\dots,m]$ at training time. Based on the definition of the attractor $A_{g_{\theta}}$ it can be proven that
$\lim_{m\rightarrow\infty}\bigcap_{i=1..N}\{g_{\theta}^{0}(\mathbf{\tilde{u}}^{t}),g_{\theta}^{1}(\mathbf{\tilde{u}}^{t}),...,g_{\theta}^{m}(\mathbf{\tilde{u}}^{t})\}\supset A_{g_{\theta}},$ (10)
@@
for sufficiently many $\mathbf{\tilde{u}}^{t}$ in the basin $\mathcal{B}(A_{g_{\theta}})$. Details of the proof are provided in A. In contrast to one-step training without unrolling, enabling it thus exposes the inference attractor $A_{g_{\theta}}$ at training time for sufficiently large datasets and horizons $m$. While the horizon $m$ is practically never chosen long enough to expose $A_{g_{\theta}}$ with a single initial condition, a sufficiently large dataset ensures that $A_{g_{\theta}}$ is approached from multiple angles and observed states from this set converge to the attractor of $g_{\theta}$ for large enough $m$. Precisely the difference between the observed training set (sampled from $A_{\mathcal{F}}$) and the inference attractor $A_{g_{\theta}}$ represents a form of data shift that can impede inference accuracy. The dark red curve in Figure 1 shows how this shift is thus reduced by choosing larger $m$.
Figure 2: Left: Illustration of unrolling for a horizon $m=2$, with numerical solver $\mathcal{S}$ and neural network $f_{\theta}$, visualized for correction and prediction chains; importantly, it highlights differences in the gradient flow shown with dotted lines in the backward pass for non-differentiable (blue, no-gradient (NOG)) and differentiable (orange, with-gradient (WIG)) setups: note how the gradients do not flow through a time step of $\mathcal{S}$ for NOG, instead they contain two shorter, disconnected backpropagation chains; Right: Contributions of networks and numerical solvers in possible prediction and correction paradigms
@@ 2.4 Training neural networks to solve PDEs:
Since the training procedure is largely similar for correction and prediction, we will use the generalized $g_{\theta}$ for the following analysis. We will put special emphasis on the differences between correction and prediction setups when they arise. The general training procedure is given as
$\theta^{*}=\mathrm{arg}\min_{\theta}\bigg{[}\sum_{t=1}^{N-m}\sum_{s={1}}^{m}\mathcal{L}_{2}\big{(}\tilde{\mathbf{u}}^{t+s},g_{\theta}^{s}(\tilde{\mathbf{u}}^{t})\big{)}\bigg{]}.$ (11)
For one learning iteration, a loss value is accumulated over an unrolled trajectory. We can write this total loss over the unrolled trajectory as
$\mathcal{L}=\sum_{s=1}^{m}\mathcal{L}_{2}(\tilde{\mathbf{u}}^{t+s},g^{s}_{\theta}(\tilde{\mathbf{u}}^{t}))=\sum_{s=1}^{m}\mathcal{L}^{s},$ (12)
where $\mathcal{L}^{s}$ represents the loss evaluated after a step. The simplest realization of the training procedure (11) utilizes the forward process with $m=1$, i.e. no unrolling. We refer to this variant without unrolling as one-step (ONE) in the following. This approach is common practice in literature and the resulting training procedure is easily realized with pre-computed data. In contrast, the full backpropagation through an unrolled chain with $m>1$ requires a differentiable solver $\mathcal{S}$ for the correction setup. The differentiable setup can calculate the full optimization gradients by propagating gradients through a simulation step, as illustrated by the orange arrows in Figure 2. The gradients are thus evaluated as
@@
We refer to this fully differentiable setup as with-gradient (WIG). The loss contribution $\mathcal{L}_{\mathrm{WIG}}^{s}$ originates from the $s$th step in the unrolling horizon. Gradients associated with this loss contribution are propagated to all previous network contributions, which results in the outer sum in equation (13). The individual network contributions require backpropagation through (multiple) simulator operations $g_{\theta}$, resulting in the chain-rule product in equation (13). Refer to Figure 2 for a visualization. The gradient flow through a simulator step $g$ is caluclated as $\frac{\partial g_{\theta,c}^{s+1}}{\partial g_{\theta,c}^{s}}=\frac{\partial f_{\theta}^{s}}{\partial\mathcal{S}^{s}}\frac{\partial\mathcal{S}^{s}}{\partial g_{\theta,c}^{s}}$ for correction setups. Here it becomes apparent that gradients propagate through a simulator step $g_{\theta,c}$. The numerical solver $\mathcal{S}$ is part of this operation in correction setups and needs to be differentiable with respect to the input state, i.e. $\frac{\partial\mathcal{S}(u)}{\partial u}$ has to be available. The propagated gradient values carry information for all network contributions lying earlier in the unrolled chain. Since there is no numerical solver in the prediction setup, the gradient propagation solely depends on the network and evaluates to $\frac{\partial g_{\theta,p}^{s+1}}{\partial g_{\theta,p}^{s}}=\frac{\partial f_{\theta}^{s}}{\partial g_{\theta,p}^{s}}$.
To test the effect of differentiability, we introduce a second strategy for propagating the gradients $\frac{\partial\mathcal{L}^{s}}{\partial\theta}$. If no differentiable solver is available, optimization gradients can only propagate to the network application, not through the solver. The blue arrows in Figure 2 do not pass through a simulation step. In mathematical terms, the gradients are thus evaluated as
@@
This setup is referred to as NOG. While gradients can still flow into the parameter space $\theta$ of an operation $g_{\theta}$, they cannot pass through it. Hence, $\frac{\partial g_{\theta}^{s+1}}{\partial g_{\theta}^{s}}:=0$ holds for correction and prediction setups. Refer to Figure 2 again for a visualization. In short, a simulator step in the NOG setup is not differentiable and yields no gradients with respect to previous steps. Most existing code bases in engineering and science are not fully differentiable. Consequentially, this NOG setup is particularly relevant, as it can be implemented using existing traditional numerical solvers. It is important to note a potential limitation of NOG unrolling. It only works without implementing additional differential operations when the ground truth information is available directly at the network output. For the correction formulation used within this study from equation (6), this is trivially given since the output of the network constitutes the new solution. If equation (7) is used without a differentiable solver instead, only one option exists - Moving the ground truth information to the output of the network and employing NOG unrolling. In the example of closure modeling, this could be achieved by explicitly filtering the high-resolution ground truth to derive subgrid-scale stresses [63]. However, deriving a filter is reliant on strong assumptions for the actual filter operation. The behavior of the resulting subgrid-scale quantities depends on the choice of filter [68]. Due to these approximations, a-priori derived ground-truth values for the subgrid-scale quantities generally are not ideal targets to maximize a-posteriori performance. This leads to a mismatch between the a-priori trained networks and filtered LES equations, which can make these networks ineffective [81]. This problem is circumvented by a-posteriori training of the subgrid-scale model, which in turn requires differentiable solvers. Utilizing the non-differentiable NOG strategy for closure models is thus only achievable with the correction formulation of equation (6).
@@
Figure 2 contains a visualization including the backpropagation flow through an unrolled chain. It highlights the fundamentally different nature of accumulated Jacobian terms for WIG and NOG: the former grows quadratically in $m$, i.e., inducing three terms in the example of the figure, while NOG grows linearly and induces only two Jacobian terms. The gradient calculations for our three setups, additionally including the ONE variant, are further denoted in Table 1. Further detail of the gradient calculations in correction and prediction setups is found in B.
@@ 2.5 Gradient behavior in unrolled systems
The long backpropagation chains used in WIG are missing in NOG unrolling and thus lead to the innermost product in Equation 15. At this stage, we can integrate a property of chaotic dynamics derived by [56] in this analysis. Therein, the authors show how the Jacobians $\mathbf{J_{s}}=\frac{\partial g_{\theta}^{s}}{\partial g_{\theta}^{s-1}}$ have eigenvalues larger than 1 in the geometric mean:
$\bigg{|}\bigg{|}\prod_{b=1}^{s-B}\frac{\partial g_{\theta}^{s-b}}{\partial g_{\theta}^{s-b-1}}\bigg{|}\bigg{|}>1.$ (16)
This means that the scale of the backpropagated gradients increases with unrolling length in the WIG setup. Since the gradient flow through $g_{\theta}$ is cut in the NOG approach, this also means that more information is omitted for long unrolling horizons, which increases the gradient inaccuracy of NOG. The largest contribution cut from the NOG gradients stems from the longest chains in equation (15). These long chains scale with $m^{2}$ due to the forward and backward pass (equivalent to the maximum values of the iterants $s$ and $B$ in equation (15)). We can thus conclude that the NOG gradient inaccuracy grows with $\propto m^{2}$, while the NOG gradients only linearly depend on $m$. For increasing unrolling lengths, the error in this surrogate gradient will outgrow the gradient itself. As a consequence, this approximation of the gradient computed in the NOG system diverges from the true (WIG) gradients for increasing $m$, as shown with the blue line in Figure 1. Due to the fact that numerically covering a characteristic timescale $t_{c}$ usually requires many timesteps, the NOG gradient divergence will be fast compared to this timescale. Using the gradient approximations used in NOG as a learning signal does not accurately match the loss used in training. This hinders the network optimization, which is visualized by the gradient divergence curve in Figure 1.
Let us finally consider the full gradients of the WIG setup. We can use Theorem 2 from [56] stating
@@
for almost all points on $A_{g_{\theta}}$. The gradients of the WIG setup explode for long horizons $m$. As a direct consequence, the training of the WIG setup becomes unstable for chaotic systems when $m$ grows. In contrast to the NOG gradients discussed above, it is not an inaccuracy in the computation that makes the training become unstable, but rather this intrinsic property of the Jacobian of chaotic dynamics. As such, the gradient divergence of the WIG system does not depend on the number of discrete steps like the NOG ones did, but is rather governed by the characteristic physical timescale $t_{c}$ of the learned system. Due to the fact that one physical timescale of a PDE system is usually discretized by numerous timesteps, the NOG gradients diverge earlier compared to WIG gradients for increasing $m$. This is indicated by the orange line below the blue one in Figure 1.
It is important to note that stopping the gradient as done by the NOG approach limits some aspects of neural network design. While exact gradients are available for the network outputs with respect to the network parameters, the same is not true for the unrolled trajectory. Consequently, a sensitivity analysis with respect to individual network parameters or design choices [69, 111] would not be possible for the entire trajectory, but rather only with respect to the direct network output. Similarly, interactions between initial variables, parameters and the output like in [29, 92] is only possible in a fully differentiable setting.
2.6 Curriculum learning
As illustrated above, unrolling can have negative effects on the gradient stability. This effect can even be amplified for newly initialized networks with random weights, where erroneous outputs make even the forward process unstable. As such, it is common practice to limit the unrolling horizon in the initial phases of the training run [93, 50]. The number of unrolled steps is then increased incrementally, which we term curriculum in this study. Curriculums are necessary when training with long unrolling horizons. A simple approach is to start with one-step learning, with longer horizons following afterward, e.g., $m=[1,2,4]$.
2.7 Summary
We can summarize the hypotheses stated above as follows: Unrolling training trajectories for data-driven learning of chaotic systems can be expected to reduce the data-shift, as the observed training samples converge to the learned attractor. At the same time, long unrollings can lead to unfavorable gradients for both NOG and WIG setups. This means that intermediate ranges of unrolling horizons can exist, where the benefits of reducing the data-shift outweigh instabilities in the gradient. Based on this trade-off, our goal is to identify preferred training regimes in our experiments where models can be trained to higher accuracy without destabilizing the learning process. We also expose the unrolling horizons and quantify the expected benefits of training with NOG and WIG modalities.
@@ 3 Physical Systems and Architectures
The four physical systems shown in Figure 3 were used for our learning tests. We focus on several widely used and well-understood test-cases from the scientific machine learning community. All individual systems are parameterized to exhibit varying behavior within the dataset, and each test set contains unseen values inside and outside the range of the training data set.
Figure 3: Visualizations of our physical systems, from left to right: KS equation state over time, KOLM vorticity field, WAKE vorticity field, AERO local Mach number field visualization for different upstream Mach numbers M